# SVD-Free Low-Rank Adaptive Gradient Optimization for Large Language Models

### v1

@@ 3.3 DCT-AdamW
We propose DCT-AdamW, a standalone low-rank version of AdamW with DCT projection that has the option to use quantized error feedback (EF) (Seide et al., 2014; Karimireddy et al., 2019; Alistarh et al., 2018) and ensures momentum buffers integrate gradients from the same lower dimensional subspaces as in LDAdamW (Robert et al., 2025). In contrast to LDAdamW, which has to store two consecutive projection matrices per layer, we only have to store two sets of $r$ indices due to simplicity of our DCT approach coupled with the dynamic column selection technique. Prior work MicroAdam (Modoranu et al., 2024) quantized the error feedback down to 4-bits in the context of compressing the optimizer state using sparsity. In our context, we observed that the lowest resolution we can quantize EF to is 8-bits. We state our optimizer in Algorithm 1.
Algorithm 1 DCT-AdamW (right projection)
1:Input: $\beta_{1},\beta_{2},\epsilon,T,T_{u},r$
2:$m_{0},v_{0}\leftarrow 0_{n\times r},0_{n\times r}$
3:$\Xi_{1}\leftarrow 0_{n\times m}$diamond error feedback (EF) buffer
4:$Q\in\mathbb{R}^{m\times m}$diamond DCT matrix
5:$\mathcal{I}_{crt},\mathcal{I}_{prev}\leftarrow 0_{r},0_{r}$diamond column indices
6:for $t=\{1,2,\dots,T\}$ do
7:     $G_{t}\leftarrow\nabla_{\theta}f(\theta_{t})+\Xi_{t}$
8:     $R\leftarrow\textsc{UpdateSubspace}(G_{t})$
9:     $g_{t}\leftarrow G_{t}\cdot Q_{crt}$diamond projected gradient
10:     $\Xi_{t}\leftarrow G_{t}-g_{t}\cdot Q_{crt}^{\top}$diamond update EF
11:     $m_{t}\leftarrow\beta_{1}\cdot m_{t-1}\cdot R+(1-\beta_{1})g_{t}$
12:     $v_{t}\leftarrow\beta_{2}|v_{t-1}\cdot R|+(1-\beta_{2})g_{t}^{2}$
14:end for
Algorithm 2 Update Subspace procedure
1:procedure UpdateSubspace($G$)
2:     $R\leftarrow I_{r\times r}$diamond switching matrix between
3:     consecutive low-rank spaces
4:     if $t>1$ then
5:         $\mathcal{I}_{prev}\leftarrow\mathcal{I}_{crt}$
6:     end if
7:     if $(t=1)\lor(t\mod T_{u}=0)$ then
8:         re-rank the columns and
9:         $\mathcal{I}_{crt}\leftarrow\textsc{RankCols}(GQ,r)$
10:         update the switching matrix
11:         $R\leftarrow Q_{prev}^{\top}\cdot Q_{crt}$
12:     end if
13:     return $R$
14:end procedure
The sets $\mathcal{I}_{prev/crt}$ hold the indices of the $r$ columns for the previous/current projections and $Q_{prev/crt}$ contain the columns from $Q$ specified by the these two sets. $T_{u}$ represents the subspace update interval (set to 200 for GaLore and to $1$ for LDAdam). The procedure RankCols ranks the columns of $S=G_{t}\cdot Q$ as explained in Section 3.2. In order to make sure the momentum buffers integrate gradients from the same subspaces, we need to rotate $m_{t}$ and $v_{t}$ using a rotation matrix $R$ that first projects the momentum to the full-dimensional space using the previous projector and then projects it back to the current lower-dimensional subspace. We can perform this rotation directly in the $r$-dimensional space by multiplying the two projection matrices, resulting in $R\in\mathbb{R}^{r\times r}$. Rotating $v_{t}$ might introduce negative values and we apply the absolute value function to force the non-negativity of $v_{t}$. The low-rank gradient $g_{t}$ is computed by multiplying the full-rank gradient $G_{t}$ with $Q_{crt}$ and the full-rank update is obtained multipying $u_{t}$ by $Q_{crt}^{\top}$.
4 Memory Savings and Running Time
In this section we provide details about running time and memory savings after replacing the expensive SVD with DCT, coupled with the dynamic column selection approach.
4.1 Memory Savings
For simplicity, we compute the memory usage for a generic model with $L$ layers, each of size $n\times n$ for a fixed rank $r$ and provide concrete values for the Llama-2 7B model, for which $n=4096$.
SVD. To perform the low-rank decomposition for a gradient matrix $G\in\mathbb{R}^{n\times n}$, we use the SVD decomposition $G=U\Sigma V^{\top}$ and use the first $r$ columns of $U$ as a projection matrix $P=U_{:,:r}$. Storing one projection matrix $P$ for each of the $L$ layers in the neural network leads to storing additional $Lnr$ elements only for the projection matrices.
DCT. To perform low-rank decomposition using DCT, we have to store the DCT matrix $Q\in\mathbb{R}^{n\times n}$ that will be the same for all layers and $r$ additional indices at each layer, which will be used to extract the corresponding columns. Finally, the total number of elements to be stored is $n^{2}+Lr$. Usually, $n$ is the hidden dimension of the model $d_{\rm model}$, which the memory usage directly depends on.
Practical Memory Usages. We compute the memory usage for a Llama-2 7B model to understand the practical implications of using a constant orthogonal matrix compared to SVD-based projections. Table 2 shows the memory usage of our approach is not sensitive to rank, since we only store $r$ integers instead of individual projection matrices with $r$ columns. Instead, the memory usage of SVD depends on both hidden dimension of the model $n=d_{\rm model}$ and the rank, since $r$ columns have to be stored. This translates to memory savings for our approach for LLMs. In DDP settings, we store one projection matrix for each GPU. However, when the model size becomes too large, GPUs can store different columns of the matrix to prevent materializing the full projection matrix on each GPU.
Table 2: Practical memory usage for low-rank projection matrices for Llama-2 7B.
SVD DCT
$r=32$ $r=256$ $r=512$ $r=32$ $r=256$ $r=512$
bfloat16 56 MiB 448 MiB 896 MiB 32.03 MiB 32.22 MiB 32.44 MiB
float32 112 MiB 896 MiB 1792 MiB 64.03 MiB 64.22 MiB 64.44 MiB
4.2 Running Time
Once the projection matrices $P$ are computed and stored, the computational complexity of the up/down-projection operations is the same as for SVD, as each involves a matrix multiplication between the projection matrix (or its transpose) and the full/low-rank gradient. However, since our projection matrices are based on the Discrete Cosine Transform (DCT), we can leverage fast Fourier transform (FFT) algorithms to reduce the complexity of the up/down-projection operations from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$. It is important to note that our method for computing the projection matrices is fundamentally different from existing approaches. In this work, we primarily focus on comparing the runtime of our method with that of SVD-based approaches.
For a generic matrix $G\in\mathbb{R}^{n\times n}$, the computational complexity of SVD is $\mathcal{O}(n^{3})$. This complexity has to be incurred for each layer once at $T_{u}$ steps because this procedure is too expensive to be run with higher frequency. On the other hand, our approach computes $S=GQ$, ranks the columns in $S$ by computing $L_{1}$-norm per column and selects the top-$r$ columns to form the projection matrix $P$ to be used for up/down-projection. At steps where we recompute the projection, we can avoid an additional matrix multiplication between $P$ and the full rank gradient matrix $G$ by returning the largest $r$ columns of $S$ as the low-rank gradient. Overall, our approach has a lower computational overhead compared to the SVD approach, which can also be seen in the running time row of results in experimental section.
@@ 5.1 Optimality of norm-based ranking procedure
We can extend the above argument for any $p$-norm over vectorized matrices, for which the Frobenius norm is the special case of $p=2$. Then, we have
@@
=$\overset{(b)}{=}\textstyle\sum_{i=r+1}^{n}\left\|q_{i}\right\|_{p}\left\|q_{i}^{\top}G\right\|_{p}\overset{(c)}{\leq}\max(1,n^{\frac{1}{p}-\frac{1}{2}})\sum_{i=r+1}^{n}\left\|q_{i}^{\top}G\right\|_{p},$
@@ 5.2 DCT as linear approximation of the gradient eigenbasis
Notice the analogy between this matrix decomposition and the Taylor expansion for functions. Since the space of matrices is finite-dimensional, the signal-matrix can be recovered using finitely many “variable” $D$’s, in contrast to the infinite series required for function expansions. Keeping this analogy in mind, just as loss functions are linearly approximated at each iteration in first-order optimization algorithms, we consider a “linear” approximation of the decomposition (2) by including only a single variable $D$, i.e., $GG^{\top}\approx FD_{1}F^{*}$.

### v2

@@ 2.4 DCT-AdamW
We propose DCT-AdamW, a standalone low-rank version of AdamW with DCT-based projection that has the option to use quantized error feedback (EF) (Seide et al., 2014; Karimireddy et al., 2019; Alistarh et al., 2018) and ensures momentum buffers integrate gradients from the same lower dimensional subspaces in a similar way as in LDAdamW (Robert et al., 2025). In contrast to LDAdamW, which has to store two consecutive projection matrices per layer, we only have to store two sets of $r$ indices per layer. Prior work MicroAdam (Modoranu et al., 2024) quantized the error feedback down to 4-bits in the context of compressing the optimizer state using sparsity. In our setup, the lowest resolution we can quantize EF to is 8-bits without degrading the optimizer performance. or the space constraints reasons, we present the details of our DCT-AdamW optimizer in Appendix E.
3 Experiments
In this section we present our numerical results. Our main goal is to show that our DCT-based dynamic column selection approach at least recovers the performance of the original algorithms which we integrate it in. Specifically, we evaluate Trion against Dion, where we directly compare the DCT-based projection followed by Newton-Schulz iteration with Power-Iteration. In addition, we replace the SVD-based projection in FRUGAL and FIRA optimizers with our DCT approach. In the end, we compare LDAdamW with our standalone DCT-AdamW optimizer. In our pretraining experiments we compare training/validation perplexity and for fine-tuning we compare the evaluation accuracy, as well as memory usage and running time for both scenarios. In the remaining part of the paper we focus on presenting our pretraining (PT) results and the fine-tuning (FT) results are presented in Appendix H.
We train from scratch models from the Llama family with 350M, 800M and 1.3B parameters using Chinchilla-optimal token counts (20 tokens per parameter) from the C4 (Raffel et al., 2020) dataset and sequence length 512. All PT experiments are run in Distributed Data Parallel (DDP) settings on 8x H100 Nvidia GPUs using batch size 512 with local batch size 64 per GPU (unless otherwise specified explicitly).
PT with Trion. Our purpose is to show that our DCT-based dynamic column selection approach can successfully replace the QR-based Power-Iteration procedure in Dion and at least recovers the performance for the best hyper-parameters reported by the original work. In Table 1, we present our results, which are obtained using the optimal learning rate $\eta=0.01$ (as reported by the original Dion paper) and weight decay $\lambda=0.01$. We report the average values acros $3$ seeds for train/validation loss and perplexity, maximum allocated memory (read directly from the GPU) and the lowest running time across all runs. We experiment with ranks 128, 256 and 512 and the rank-to-dimension ratio is $r/d\in\{1/16,1/8,1/4,1/2\}$, equivalent to $6.25\%,12.5\%,25\%$ and $50\%$ of the full rank $d$, where $d$ is model’s embedding dimension.
Rank $r$ Metric 350M ($d=1024$) 800M ($d=2048$) 1.3B ($d=2048$)
Trion Dion Trion Dion Trion Dion
128 $r/d$ $1/8$ $1/16$ $1/16$
Train Loss 2.726 2.764 2.532 2.555 2.475 2.503
Train PPL 15.29 15.89 12.59 12.89 11.90 12.24
Val Loss 2.736 2.773 2.505 2.527 2.422 2.448
Val PPL 15.43 16.00 12.25 12.52 11.27 11.56
Memory (GB) 42.42 45.56 67.45 71.64 63.62 68.57
Runtime 1h 53m 1h 58m 8h 5m 8h 24m 19h 13m 19h 42m
256 $r/d$ $1/4$ $1/8$ $1/8$
Train Loss 2.715 2.741 2.533 2.546 2.476 2.492
Train PPL 15.11 15.51 12.61 12.78 11.92 12.11
Val Loss 2.727 2.750 2.503 2.519 2.423 2.440
Val PPL 15.30 15.64 12.22 12.42 11.28 11.47
Memory (GB) 42.42 45.59 67.45 71.75 63.62 68.58
Runtime 1h 53m 2h 3m 8h 3m 8h 36m 19h 13m 20h 6m
512 $r/d$ $1/2$ $1/4$ $1/4$
Train Loss 2.708 2.718 2.528 2.535 2.470 2.481
Train PPL 15.01 15.16 12.54 12.63 11.84 11.97
Val Loss 2.718 2.727 2.499 2.509 2.420 2.430
Val PPL 15.15 15.29 12.18 12.30 11.25 11.36
Memory (GB) 42.42 45.99 67.45 71.81 63.62 68.73
Runtime 1h 54m 2h 14m 8h 3m 8h 48m 19h 13m 20h 44m
Table 1: Perplexity, Memory, and Running Time Comparison for Trion and Dion. $d$ stands for the embedding dimensionality of the model. The 1.3B model was trained with local batch size 32
Performance In Table 1, we show Trion consistently achieves lower training and validation loss, which also translates to lower perplexities. At the top row of Figure 3 we show the training loss curve of Trion is lower than Dion over time. This is an indication that our DCT-based dynamic column selection algorithm, followed by orthogonalization via Newton-Schulz can successfully replace the Power-Iteration procedure in Dion.
Memory Usage. In Table 1, we show Trion has a lower memory requirement across all experiments since it allocates only one DCT matrix per GPU of size $d\times d$, from which we select $r$ columns using the indices of the most significant columns. In contrast, Dion stores a projection matrix for each layer. This design difference translates to around $10\%$ lower memory footprint for Trion compared to Dion.
Runtime. We would like to emphasize that running time of Trion does not depend on rank, as can be seen from the reported runtime across different ranks in Table 1: Trion achieves nearly constant runtime for each model size across all ranks, while the runtime of Dion clearly depends on the rank because of the QR-decomposition. This represents an advantage of Trion for much larger scales compared to Dion. At the bottom row of Figure 3 we present the wall clock time for the two optimizers for rank $256$ across all three models we tested on. Concretely, for a fixed running time budget, Trion consistently achieves lower training loss than Dion. Trion is faster than Dion by $2.5-4.5\%$ for rank $128$, $4.5-9\%$ for rank $256$ and $8-18\%$ for rank $512$. It is important to mention that the embedding size $d$ of the models we used in our work is too small to see a significant difference in the runtime between Makhoul’s algorithm and matmul when computing the column similarities. However, our benchmark presented in Appendix D shows the advantage of using Makhoul’s algorithm at scale compared to matmul.
[…]
PT with LDAdamW. We train Llama-800M on 80B tokens (100 tokens per parameter) from C4 using LDAdam and DCT-AdamW described in Algorithm 2. In particular for our approach, we use EF quantized to 8-bits and some further optimizations from the ZeRO-redundancy optimizer (Rajbhandari et al., 2020), where one layer is updated on a single GPU and then it is broadcasted to the other GPUs. This way, we obtain lower memory usage by replacing redundant operations in the optimizer with communication, since the GPUs receiving the updated layer parameters do not allocate any state, at the cost of slightly increased running time due to more communication. In Figure 2 we present the training loss curves for full-rank AdamW (for reference), LDAdamW and DCT-AdamW and we are interested in directly comparing the last two. We observe that DCT-AdamW has lower training loss than LDAdamW, which also translates to lower perplexities in Table 2. Due to relatively high rank, the memory usage of LDAdamW is close to the memory usage of AdamW because it stores two projection matrices to be able to rotate the momentum buffers. In contrast, DCT-AdamW stores only two sets of $r$ indices instead of storing the actual projection matrices, which drastically reduces the memory usage, coupled with the ZeRO-redundancy trick. In terms of running time, DCT-AdamW is faster than LDAdamW by 10h 7m ($\approx 25.75\%)$ and slower than AdamW by 1h 55m ($\approx 5\%)$.
Figure 1: Projection errors for Dion and Trion for a few layers from the first transformer block on Llama-30M ($d=640,r=128$).
AdamW LDAdamW DCT-AdamW
Train PPL 12.88 15.10 14.95
Val. PPL 11.73 13.91 13.69
Mem. (GiB) 73.72 72.10 57.82
Time 1d 13h 22m 2d 1h 24m 1d 15h 17m
(a) Llama-350M
(b) Llama-800M
(c) Llama-1.3B
Figure 3: Training Loss across iterations (top row) and wall clock time (bottom row) for three Llama models trained with Dion and Trion optimizers on C4 using 20 tokens per parameter and rank $r=256$. For a clearer visualization, the training loss curve was smoothed using a moving average with a window size of 200 and we zoom-in on the last 5000 training steps.
PT with FRUGAL/FIRA. We replace the SVD projection with our DCT approach and test it on Llama-800M model using 16B tokens from the C4. For the space constraints reasons, we present the results in Appendix G.
@@ 4.1 Optimality of norm-based ranking procedure
We can extend the above argument for any $p$-norm over vectorized matrices, for which the Frobenius norm is the special case of $p=2$. Then, we have:
@@
$\overset{(b)}{=}\textstyle\sum_{i=r+1}^{n}\left\|q_{i}\right\|_{p}\left\|q_{i}^{\top}G\right\|_{p}\overset{(c)}{\leq}\max(1,n^{\frac{1}{p}-\frac{1}{2}})\sum_{i=r+1}^{n}\left\|q_{i}^{\top}G\right\|_{p},$
@@ 4.2 DCT as linear approximation of the gradient eigenbasis
Notice the analogy between this matrix decomposition and the Taylor expansion for functions. Since the space of matrices is finite-dimensional, the signal-matrix can be recovered using finitely many “variable” $D$’s, in contrast to the infinite series required for function expansions. Keeping this analogy in mind, just as loss functions are linearly approximated at each iteration in first-order optimization algorithms, we consider a “linear” approximation of the decomposition equation 2 by including only a single variable $D$, i.e., $GG^{\top}\approx FD_{1}F^{*}$.